# Die Hard: The On-Off-Cycle of Galaxies on the Star Formation Main Sequence

### v1

@@ 1 Introduction
The transformation of diffuse gas into stars is a process that spans from large scale structure formation (e.g. van de Weygaert & Bond 2008) down to giant molecular clouds that host star forming clumps (e.g. McKee & Ostriker 2007). As baryonic physics is subject to a chaotic interplay of gravitation and hydrodynamics with cooling, heating and dust formation as well as feedback from stars, supernovae and active galactic nuclei (AGN), it is remarkable that structure and star formation in galaxies produces characteristic relations and evolutionary pathways. In a color-magnitude diagram, active galaxies form a “blue cloud” separated from the “red sequence” (e.g. Strateva et al. 2001; Balogh et al. 2004) with fewer galaxies occupying the “green valley” in between (Wyder et al. 2007). Additionally, active galaxies are observed to follow a relation between their stellar masses and star formation rates, which is called the “star formation main sequence” (e.g. Speagle et al. 2014; Pearson et al. 2018; Leslie et al. 2020; Popesso et al. 2023). Consistent findings of a declining star formation main sequence toward lower redshifts show that the cosmological conditions for conversion of gas into stars change with redshift (e.g. Teklu et al. 2023). This can also be seen from studies of the star formation rate density (SFRD), which evolves with a universal peak at $z\approx 2$ and lower total SFRD at both higher and lower redshifts (Madau & Dickinson 2014). Their results for the SFRD evolution are consistent with a general decline of specific star formation rates ($\mathrm{sSFR}=\mathrm{SFR}/\mathrm{M}_{*}$) in galaxies at later times, which corresponds with a decline of the overall star formation main sequence. However, the different mass- and redshift-dependent descriptions of the main sequence (e.g. Pearson et al. 2018; Santini et al. 2017; Speagle et al. 2014) show that the physics behind the scaling relation are not yet fully understood.
@@
Lin et al. (2024) used ALMA and MaNGA observations and found “green-valley” galaxies that displayed low star-formation rates despite having similar amounts of dense gas as main-sequence galaxies. Even with constant and smooth accretion, internal processes of a galaxy are expected to lead to a scatter of $\Delta\mathrm{MS}=\mathrm{log}(\mathrm{SFR}/\mathrm{SFR_{\mathrm{MS}}})\approx 0.3$ dex around the main sequence at different redshifts (e.g. Noeske et al. 2007; Whitaker et al. 2012; Speagle et al. 2014; Pearson et al. 2018). Tacchella et al. (2016) tested the scatter at $z>1$ using results from the Vela zoom-in simulation suite and attribute this to a cycle of compaction, depletion and accretion in galaxies until gas inflow from the cosmic web comes to an end. Mergers provide an additional mode of increased star formation activity, either as a consequence of a wet-merger-induced starburst or through enhanced gas compaction via interaction with a galaxy (Gómez-Guijarro et al. 2022). In our current picture of galaxy formation, it is usually assumed that galaxies evolve along the main sequence until they are quenched by one of the commonly discussed processes, such as galaxy mergers (e.g. Kormendy & Ho 2013; Ellison et al. 2022) or halo mass quenching (Birnboim & Dekel 2003; Dekel & Birnboim 2008). However, cases of rejuvenation have been discussed lately, as evidence for this is seen at high redshifts (Remus & Kimmig 2025; Trussler et al. 2025) but also at low redshifts (e.g. Yi et al. 2005; Chauke et al. 2019). Recently, models also predict that rejuvenation might be a common occurrence in galaxy formation (e.g. Zhang et al. 2023; Tanaka et al. 2024). However, as of yet it is unclear how common this process really is and how often a galaxy simply evolves along the main sequence, as commonly assumed.
@@
The principle of mass conservation in its application as the “Bathtub” gas regulator model for smooth accretion (e.g. Bouché et al. 2010; Lilly et al. 2013; Burkert 2017) provides a simple framework to test our understanding of the interplay of gas flows and star formation. This interplay is governed by the depletion time $t_{\mathrm{depl}}\equiv M_{\mathrm{gas}}/\mathrm{SFR}\sim 1\ \mathrm{Gyr}$ (Tacconi et al. 2020), where processes like recycling of gas in a “halo fountain” (Oppenheimer & Davé 2008; Tumlinson et al. 2017) and cooling become important. In order to identify typical star formation histories and quantify their prevalence, we study and classify long-term trends of galaxies in the Magneticum Pathfinder cosmological simulations111www.magneticum.org. This paper is structured as follows: In section 2, we describe the simulation used for this work. Then, we describe in section 3 the method employed to identify moments of quenching and rejuvenation as well as how we use them to classify the formation histories of the galaxies in our sample. We then show how these classes differ at $z\approx 0$ with respect to their mean ages, metallicities and mass distributions in section 4. Finally, we investigate causes of the different characteristic formation classes by measuring gas flow rates, gas flow geometries and environment in sections  5 and  6.
@@ 2 The Magneticum Simulations
We perform our analysis using Box4 uhr of the Magneticum Pathfinder cosmological hydrodynamical simulation suite (Teklu et al. 2015). These simulations are based on the cosmological parameters from the WMAP7 results according to Komatsu et al. (2011) with a mass density of $\Omega_{\mathrm{m}}=0.272$, a baryon fraction of $f_{\mathrm{baryon}}=0.168$, a cosmological constant of $\Lambda=0.728$ and a dimensionless Hubble constant of $h=0.704$. The simulations were run with a modified version of the SPH code GADGET-2 (Springel et al. 2001b; Springel 2005) with updates regarding the SPH kernel and treatment of viscosity (Dolag et al. 2005; Beck et al. 2016; Donnert et al. 2013).
@@
The baryonic physics includes gas cooling, star formation and stellar winds as described by Springel & Hernquist (2003). Accordingly, gas particles include contributions to the hot and cold phases with star formation based on the latter. Exchange between the two phases is implemented through radiative cooling on one hand with rates based on cooling tables from Wiersma et al. (2009). Their results are based on calculations using the code cloudy (Ferland et al. 1998) while considering contributions from the eleven elements H, He, C, N, O, Ne, Mg, Si, S, Ca and Fe, which are traced individually within the simulation for all stellar and gas particles, with exposition to the cosmic microwave background radiation and the UV/X-ray background radiation from galaxies and quasars according to the model by Haardt & Madau (2001).
@@
Star formation, supernovae (SNe) and chemical enrichment are implemented as a subresolution model according to Tornatore et al. (2003, 2007) with initial stellar population distribution according to Chabrier (2003). The SN feedback is implemented partly as heating and partly as galactic winds with $v_{\mathrm{wind}}=350$ km/s triggered from a 10%-fraction of SNe II ($10^{51}$ erg). Black hole growth and AGN feedback are modeled as described by Springel et al. (2005); Di Matteo et al. (2005) with modifications covered by Fabjan et al. (2010) and Hirschmann et al. (2014). AGN luminosities are derived from the BH accretion rate as $L_{\mathrm{r}}=\epsilon_{\mathrm{r}}\ \dot{M}_{\bullet}\ c^{2}$, with a radiative efficiency of $\epsilon_{\mathrm{r}}=0.1$ for a non-rapidly spinning black hole (Shakura & Sunyaev 1973). Of the total radiated energy, $\epsilon_{\mathrm{f}}=10\%$ is thermally coupled to the surrounding gas, leading to a feedback energy rate of $\dot{E}_{\mathrm{f}}=\epsilon_{\mathrm{r}}\ \epsilon_{\mathrm{f}}\ \dot{M}_{\bullet}\ c^{2}$. Thermal conduction in galaxy clusters is implemented as described by Dolag et al. (2004) but with a value of $1/20$ (instead of $1/3$) of the Spitzer value (Spitzer 1962), since anisotropic conduction effects create a net heat transport corresponding to a Spitzer value of one percent, while values of more than ten percent lead to oversmoothed temperature profiles, as shown by Arth et al. (2014).
With a box size of $48\ \mathrm{h}^{-1}\mathrm{Mpc}$, the simulation volume contains 1318 central galaxies with a final $z=0$ stellar mass of $M_{*}\geq 1\times 10^{10}M_{\odot}$. We choose central galaxies with this stellar mass threshold in order to reliably trace galaxies, including their kinematics, with sufficient particle resolution of $N_{*}>5000$. This is a consequence of the resolution given by particles masses of $m_{\mathrm{dm}}=3.6\times 10^{7}\mathrm{h}^{-1}M_{\odot}$ and $m_{\mathrm{gas}}=7.3\times 10^{6}\mathrm{h}^{-1}M_{\odot}$ and softening lengths of $\epsilon_{\mathrm{dm/gas}}=1.4\mathrm{h}^{-1}\mathrm{kpc}$ and $\epsilon_{\mathrm{*}}=0.7\mathrm{h}^{-1}\mathrm{kpc}$. Each gas particle can spawn a star particle up to four times resulting in lower masses for stellar particles of $\mathrm{m}_{*}\approx 1.8\cdot 10^{6}\mathrm{h}^{-1}M_{\odot}$. The fact that gas particles are not fully consumed upon forming a star is a unique advantage of the Magneticum Pathfinder simulations when studying gas flows, as it allows star forming gas, which typically sits deep in the potential and thus experiences high chemical enrichment from the surrounding stars, to both form a star at its current metallicity, then be further enriched and ejected by the stellar feedback, before being reaccreted in cycles to form another star later.
@@
The simulation has been shown to reproduce several critical galaxy properties in previous works, for example with respect to scaling relations (e.g. Teklu et al. 2015, 2017; Remus & Forbes 2022), the kinematical properties of galaxies (e.g., Schulze et al. 2018; van de Sande et al. 2019; Valenzuela et al. 2024), and the star formation properties of galaxies (e.g. Kudritzki et al. 2021; Teklu et al. 2023). It has also been shown to already capture galaxy properties at high redshifts, which is of particular difficulty (e.g. Lustig et al. 2023; Remus & Kimmig 2025; Kimmig et al. 2025), but especially relevant in the context of this study where we trace galaxy properties through cosmic time.
@@
We traced a set of galaxies by following the most massive progenitors along halo merger trees of SUBFIND-identified galaxies with a baryonic adaptation by Dolag et al. (2009) of the work by Springel et al. (2001a). We disregard galaxies with an abrupt termination of their merger tree at $z<1$, as well as galaxies that do not enter the star forming main sequence at $z>1$ and by that reduce our final sample from 1318 to 1282 galaxies.
@@
Table 1 in subsection 3.1 shows that the latter criterion of a period of star formation on the main sequence impacts the total number of galaxies considered.
@@
Figure 1: The star formation main sequence at $z\approx 0$. The left- and right-hand panels show the results with respect to two selection criteria for Magneticum galaxies. By imposing a threshold of $\mathrm{sSFR}>0.2/t_{\mathrm{Hub}}$ (left) as used by Carnall et al. (2020) (dashed black line), we find a slightly higher linear fit to the logarithmic values (dashed blue line) compared to selecting for all kinematic disks with $\mathrm{SFR}>0\ \mathrm{M}_{\odot}/\mathrm{yr}$ (right). Results from observations are displayed in magenta, simulations are shown in blue colors. The solid blue line and shaded region show the running median and 50th-percentile region. The circles represent the galaxies from the Magneticum Pathfinder simulations colored according to their morphology as defined by the b-value (Teklu et al. 2015). Gray circles have been disregarded by the respective criterion. The magenta solid line and shaded region highlights a range of ±$\pm 0.5\ \mathrm{dex}$ around the main sequence fit by Pearson et al. (2018), which we used in this work to distinguish between on and off the main sequence. The other magenta lines are additional observational results as dashed (Speagle et al. 2014), dash-dotted (Leslie et al. 2020) and dotted (Popesso et al. 2023). The dashed cyan line marks the results from Illustris TNG by Donnari et al. (2019a).
@@
This is especially evident when comparing the total number of galaxies classified with respect to the main sequence by Speagle et al. (2014), for which only 1014 galaxies are among the formation classes discussed in this work.
@@ 3 The Main Sequence and Star Formation History Classes
The relation of galaxy stellar mass and star formation rate has long been known to establish a fundamental difference between two groups of galaxies, with those galaxies that are star forming living on a relatively tight star forming main sequence, and those that have quenched their star formation lying within a more diffuse cloud below (e.g. Speagle et al. 2014; Pearson et al. 2018; Popesso et al. 2023; Leslie et al. 2020; Brinchmann et al. 2004). This main sequence of galaxies evolves with time, reaching higher star formation rates at a given stellar mass at higher redshifts than at lower redshifts, demonstrating an overall decline in star formation to lower redshifts (e.g., Speagle et al. 2014; Pearson et al. 2018). While this global picture is well established, the observed main sequence relations strongly differ between different surveys already at $z=0$, as can be seen from Fig. 1, where the different observed relations are shown as pink lines.
@@
We start by testing whether the simulation can reproduce the observed main sequences. Fig. 1 shows in both panels the star forming main sequence found for the Magneticum simulations, in comparison to the results reported from different observations and the IllustrisTNG simulations (Donnari et al. 2019b, dashed cyan line). The left panel shows all galaxies colored according to their morphology as traced by the b-value (Teklu et al. 2015), with blue colors indicating disk galaxies and red colors indicating spheroidals. Yellow colors represent intermediate galaxies. The main sequence from the simulations here is calculated from all galaxies that have specific star formation rates larger than $\mathrm{sSFR}>0.2/t_{\mathrm{Hub}}$, the threshold criterion defined by Carnall et al. (2020), similar to that by Franx et al. (2008) commonly used in the literature to split quiescent from star forming galaxies and included in the figure as black dashed line. The resulting main sequence from Magneticum using this definition is marked as a blue solid line, with the 50th-percentile scatter shown as a blue shaded area. The main sequence found for the Magneticum simulations using this definition mirrors the results reported for the IllustrisTNG simulations by Donnari et al. (2019b). It well agrees with the observational relations, though we note the existence of some scatter, such that the observations by Pearson et al. (2018) tend to higher SFRs at lower stellar masses compared to Magneticum, while Leslie et al. (2020) and Popesso et al. (2023) find lower SFRs for the more massive galaxies at $z=0$.
@@
The right panel of Fig. 1 shows the main sequence as obtained if only those galaxies that are morphologically classified as disk galaxies are used. As can be seen immediately, the main sequence obtained from the simulation now is overall at lower values, resembling more the shape of the relation reported by Popesso et al. (2023) or Leslie et al. (2020) before a sudden increase of star formation at $M_{*}>10^{11.5}\mathrm{M}_{\odot}$. This is due to the existence of what is commonly referred to as red spirals (e.g. Masters et al. 2010), i.e., galaxies that are morphologically spiral galaxies but their star formation has (partially) been quenched. Their presence demonstrates that the morphology of galaxies is not directly correlated to their star formation properties, even though for the large majority there is the trend of disk galaxies being star forming. This was already indicated in the left panel of Fig. 1, where the coloring of the data points revealed that several of the galaxies on the main sequence according to their star formation properties are in fact of intermediate morphology (i.e., contain for example large bulges) or even, in a few cases, spheroidals, especially at the higher mass end.

### v2

@@ 1 Introduction
The transformation of diffuse gas into stars is a process that spans from large scale structure formation (e.g. vandeweygaert:2008:bond) down to giant molecular clouds that host star forming clumps (e.g. mckee:2007:ostriker). As baryonic physics is subject to a chaotic interplay of gravitation and hydrodynamics with cooling, heating and dust formation as well as feedback from stars, supernovae and active galactic nuclei (AGN), it is remarkable that structure and star formation in galaxies produces characteristic relations and evolutionary pathways. In a color-magnitude diagram, active galaxies form a “blue cloud” separated from the “red sequence” (e.g. strateva:2001; balogh:2004) with fewer galaxies occupying the “green valley” in between (wyder:2007). Additionally, active galaxies are observed to follow a relation between their stellar masses and star formation rates, which is called the “star formation main sequence” (SFMS) (e.g. speagle:2014; pearson:2018; leslie:2020; popesso:2023). Consistent findings of a declining star formation main sequence toward lower redshifts show that the cosmological conditions for conversion of gas into stars change with redshift (e.g. teklu:2023). This can also be seen from studies of the star formation rate density (SFRD), which evolves with a universal peak at $z\approx 2$ and lower total SFRD at both higher and lower redshifts (madau:2014:dickinson). Their results for the SFRD evolution are consistent with a general decline of specific star formation rates ($\mathrm{sSFR}=\mathrm{SFR}/\mathrm{M}_{*}$) in galaxies at later times, which corresponds with a decline of the overall star formation main sequence. However, the different mass- and redshift-dependent descriptions of the main sequence (e.g. pearson:2018; santini:2017; speagle:2014) show that the physics behind the scaling relation are not yet fully understood.
@@
lin:2024 used ALMA and MaNGA observations and found “green-valley” galaxies that displayed low star-formation rates despite having similar amounts of dense gas as main-sequence galaxies. Even with constant and smooth accretion, internal processes of a galaxy are expected to lead to a scatter of $\Delta\mathrm{MS}=\mathrm{log}(\mathrm{SFR}/\mathrm{SFR_{\mathrm{MS}}})\approx 0.3$ dex around the main sequence at different redshifts (e.g. noeske:2007a; whitaker:2012; speagle:2014; pearson:2018). tacchella:2016 tested the scatter at $z>1$ using results from the Vela zoom-in simulation suite and attribute this to a cycle of compaction, depletion and accretion in galaxies until gas inflow from the cosmic web comes to an end. Mergers provide an additional mode of increased star formation activity, either as a consequence of a wet-merger-induced starburst or through enhanced gas compaction via interaction with a galaxy (gomez-guijarro:2022). In our current picture of galaxy formation, it is usually assumed that galaxies evolve along the main sequence until they are quenched by one of the commonly discussed processes, such as galaxy mergers (e.g. kormendy:2013:ho; ellison:2022) or halo mass quenching (birnboim:2003:dekel; dekel:2008:birnboim). However, cases of rejuvenation have been discussed lately, as evidence for this is seen at high redshifts (remus:2023; trussler:2025) but also at low redshifts (e.g. yi:2005; chauke:2019). Recently, models also predict that rejuvenation might be a common occurrence in galaxy formation (e.g. zhang:2023; tanaka:2024). However, as of yet it is unclear how common this process really is and how often a galaxy simply evolves along the main sequence, as commonly assumed.
@@
The principle of mass conservation in its application as the “Bathtub” gas regulator model for smooth accretion (e.g. bouche:2010; lilly:2013; burkert:2017) provides a simple framework to test our understanding of the interplay of gas flows and star formation. This interplay is governed by the depletion time $t_{\mathrm{depl}}\equiv M_{\mathrm{gas}}/\mathrm{SFR}\sim 1\ \mathrm{Gyr}$ (tacconi:2020), where processes like recycling of gas in a “halo fountain” (oppenheimer:2008:dave; tumlinson:2017) and cooling become important. In order to identify typical star formation histories and quantify their prevalence, we study and classify long-term trends of galaxies in the Magneticum Pathfinder cosmological simulations111www.magneticum.org. This paper is structured as follows: In section 2, we describe the simulation used for this work. Then, we describe in section 3 the method employed to identify moments of quenching and rejuvenation as well as how we use them to classify the formation histories of the galaxies in our sample. We then show how these classes differ at $z\approx 0$ with respect to their mean ages, metallicities and mass distributions in section 4. Finally, we investigate causes of the different characteristic formation classes by measuring gas flow rates, gas flow geometries and environment in sections 5 and 6.
@@ 2 The Magneticum simulations
We perform our analysis using Box4 uhr of the Magneticum Pathfinder cosmological hydrodynamical simulation suite (teklu:2015). These simulations are based on the cosmological parameters from the WMAP7 results according to komatsu:2011 with a mass density of $\Omega_{\mathrm{m}}=0.272$, a baryon fraction of $f_{\mathrm{baryon}}=0.168$, a cosmological constant of $\Lambda=0.728$ and a dimensionless Hubble constant of $h=0.704$. The simulations were run with a modified version of the smoothed-particle hydrodynamics (SPH) code GADGET-2 (springel:2001b; springel:2005a) with updates regarding the SPH kernel and treatment of viscosity (dolag:2005; beck:2016; donnert:2013).
@@
The baryonic physics includes gas cooling, star formation and stellar winds as described by springel:2003:hernquist. Accordingly, gas particles include contributions to the hot and cold phases with star formation based on the latter. Exchange between the two phases is implemented through radiative cooling on one hand with rates based on cooling tables from wiersma:2009. Their results are based on calculations using the code cloudy (ferland:1998) while considering contributions from the eleven elements H, He, C, N, O, Ne, Mg, Si, S, Ca and Fe, which are traced individually within the simulation for all stellar and gas particles, with exposition to the cosmic microwave background radiation and the UV/X-ray background radiation from galaxies and quasars according to the model by haardt:2001:madau.
@@
Star formation, supernovae (SNe) and chemical enrichment are implemented as a subresolution model according to tornatore:2003; tornatore:2007; dolag:2017 with initial stellar population distribution according to chabrier:2003. The SN feedback is implemented partly as heating and partly as galactic winds with $v_{\mathrm{wind}}=350$ km/s triggered from a 10%-fraction of SNe II ($10^{51}$ erg). Black hole growth and AGN feedback are modeled as described by springel:2005b; dimatteo:2005 with modifications covered by fabjan:2010 and hirschmann:2014. AGN luminosities are derived from the BH accretion rate as $L_{\mathrm{r}}=\epsilon_{\mathrm{r}}\ \dot{M}_{\bullet}\ c^{2}$, with a radiative efficiency of $\epsilon_{\mathrm{r}}=0.1$ for a non-rapidly spinning black hole (shakura:1973:sunyaev). Of the total radiated energy, $\epsilon_{\mathrm{f}}=10\%$ is thermally coupled to the surrounding gas, leading to a feedback energy rate of $\dot{E}_{\mathrm{f}}=\epsilon_{\mathrm{r}}\ \epsilon_{\mathrm{f}}\ \dot{M}_{\bullet}\ c^{2}$. Thermal conduction in galaxy clusters is implemented as described by dolag:2004 but with a value of $1/20$ (instead of $1/3$) of the Spitzer value (spitzer:1962), since anisotropic conduction effects create a net heat transport corresponding to a Spitzer value of one percent, while values of more than ten percent lead to oversmoothed temperature profiles, as shown by arth:2014.
[…]
@@
The simulation has been shown to reproduce several critical galaxy properties in previous works, for example with respect to scaling relations across different redshifts (e.g. teklu:2015; teklu:2017; remus:2022; dolag:2025), the kinematical properties of galaxies (e.g., schulze:2018; vdsande:2019; valenzuela:2024), and the star formation properties of galaxies (e.g. kudritzki:2021; teklu:2023). It is especially noteworthy that the galactic scaling relations arise naturally while the tuning of the Magneticum Pathfinder simulations was towards hot halos of galaxy clusters. Furthermore, the simulation suite captures galaxy properties at the higher redshifts of $z\approx 4$, which is of particular difficulty (e.g. lustig:2023; remus:2023; kimmig:2025a; dolag:2025), but especially relevant in the context of this study where we trace galaxy properties through cosmic time.
@@
We traced a set of galaxies by following the most massive progenitors along halo merger trees of SUBFIND-identified galaxies with a baryonic adaptation by dolag:2009 of the work by springel:2001a. To clean the sample from rare cases of broken merger trees, we disregard galaxies with an abrupt termination of their merger tree at $z<1$, since this makes it difficult to compare with galaxies that are traced back to cosmic noon and beyond. Additionally, we remove galaxies that do not enter the star formation main sequence before $z>1$, reducing out final sample from 1318 to 1282 galaxies, which we plot in Figure 1.
@@
footnote 2 in subsection 3.1 shows that the latter criterion of a period of star formation on the main sequence impacts the total number of galaxies considered.
@@
Figure 1: The star formation main sequence at $z\approx 0$. The left- and right-hand panels show the results with respect to two selection criteria for Magneticum galaxies. By imposing a threshold of $\mathrm{sSFR}>0.2/t_{\mathrm{Hub}}$ (left) as used by carnall:2020 (dashed black line), we find a slightly higher linear fit to the logarithmic values (dashed blue line) compared to selecting for all kinematic disks with $\mathrm{SFR}>0\ \mathrm{M}_{\odot}/\mathrm{yr}$ (right). Results from observations are displayed in magenta, simulations are shown in blue colors. The solid blue line and shaded region show the running median and 50th-percentile region. The circles represent the galaxies from the Magneticum Pathfinder simulations colored according to their morphology as defined by the b-value (teklu:2015). Gray circles are ignored for the running median lines on each panel. SFR values below the range of the figure are fixed to the bottom to represent the full sample in this work. The magenta solid line and shaded region highlights a range of $\pm 0.5\ \mathrm{dex}$ around the main sequence fit by pearson:2018, which we used in this work to distinguish between on and off the main sequence. The other magenta lines are additional observational results as dashed (speagle:2014), dash-dotted (leslie:2020) and dotted (popesso:2023). The dashed cyan line marks the results from IllustrisTNG by donnari:2019err.
@@
This is especially evident when comparing the total number of galaxies classified with respect to the main sequence by speagle:2014, for which only 1014 galaxies are among the formation classes discussed in this work.
@@ 3 The main sequence and star formation history classes
The relation of galaxy stellar mass and star formation rate has long been known to establish a fundamental difference between two groups of galaxies, with those galaxies that are star forming living on a relatively tight star formation main sequence, and those that have quenched their star formation lying within a more diffuse cloud below (e.g. speagle:2014; pearson:2018; popesso:2023; leslie:2020; brinchmann:2004). This main sequence of galaxies evolves with time, reaching higher star formation rates at a given stellar mass at higher redshifts than at lower redshifts, demonstrating an overall decline in star formation to lower redshifts (e.g., speagle:2014; pearson:2018). While this global picture is well established, the observed main sequence relations strongly differ between different surveys already at $z=0$, as can be seen from Fig. 1, where the different observed relations are shown as pink lines.
@@
We start by testing whether the simulation can reproduce the observed main sequences. Fig. 1 shows in both panels the star formation main sequence found for the Magneticum simulations, in comparison to the results reported from different observations and the IllustrisTNG simulations (donnari:2019, dashed cyan line). The left panel shows all galaxies colored according to their morphology as traced by the b-value (teklu:2015), with blue colors indicating disk galaxies and red colors indicating spheroidals. Yellow colors represent intermediate galaxies. The main sequence from the simulations here is calculated from all galaxies that have specific star formation rates larger than $\mathrm{sSFR}>0.2/t_{\mathrm{Hub}}$, the threshold criterion defined by carnall:2020, while galaxies below this threshold are plotted in gray. This criterion is similar to that by franx:2008 commonly used in the literature to split quiescent from star forming galaxies and included in the figure as black dashed line. The resulting main sequence from Magneticum using this definition is marked as a blue solid line, with the 50th-percentile scatter shown as a blue shaded area. The main sequence found for the Magneticum simulations using this definition mirrors the results reported for the IllustrisTNG simulations by donnari:2019. It well agrees with the observational relations, though we note the existence of some scatter, such that the observations by pearson:2018 tend to higher SFRs at lower stellar masses compared to Magneticum, while leslie:2020 and popesso:2023 find lower SFRs for the more massive galaxies at $z=0$.
@@
The right panel of Fig. 1 shows the main sequence as obtained if only those galaxies that are morphologically classified as disk galaxies are used, with the intermediates and spheroidals plotted in gray. While this criterion does not explicitly select for star-forming galaxies, the absence of blue circles at the bottom of the panel shows that disk galaxies also have higher SFR values. As can be seen immediately, the main sequence obtained from the simulation now is overall at lower values, resembling more the shape of the relation reported by popesso:2023 or leslie:2020 before a sudden increase of star formation at $M_{*}>10^{11.5}\mathrm{M}_{\odot}$. This is due to the existence of what is commonly referred to as red spirals (e.g. masters:2010), i.e., galaxies that are morphologically spiral galaxies but their star formation has (partially) been quenched. Their presence demonstrates that the morphology of galaxies is not directly correlated to their star formation properties, even though for the large majority there is the trend of disk galaxies being star forming. This was already indicated in the left panel of Fig. 1, where the coloring of the data points revealed that several of the galaxies on the main sequence according to their star formation properties are in fact of intermediate morphology (i.e., contain for example large bulges) or even, in a few cases, spheroidals, especially at the higher mass end.